OpenStudy (anonymous):

solve 6x-5=10x-7 please help!!!!!!! Medals!!!!!!!!!!!!!! Please explain to how to get X

OpenStudy (whpalmer4):

6x-5=10x-7 You're trying to get x alone on one side of the equals sign, and an expression giving the value of x on the other side. You can add, subtract, multiply, divide, take square roots, etc. to do this, so long as you do the same thing to both sides. Let's subtract 6x from each side: 6x-5-6x = 10x-7-6x That simplifies to: -5 = 4x-7 Now we can get rid of the -7 on the right side by adding 7 to both sides: -5+7 = 4x-7+7 which simplifies to: 2=4x How do you think you find the value of x at this point?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x=.5

OpenStudy (whpalmer4):

that's correct.